OpenStudy (anonymous):

Solve for the variable in this equation. x/2-2/3=1/4

OpenStudy (thomas5267):

Is this the equation? \[ \frac{x}{2}-\frac{2}{3}=\frac{1}{4} \]

OpenStudy (anonymous):

Yes:)

OpenStudy (anonymous):

Will you help me??

OpenStudy (anonymous):

x/2 = 1/4 - 2/3 x/2 = 3/12 - 8/12 x/2 = -5/12 2* x/2 = (-5/12) * 2 x = -10/12 x = -5/6

OpenStudy (anonymous):

Do you see how I did that?

OpenStudy (thomas5267):

\[ \begin{align*} \frac{x}{2}-\frac{2}{3}&=\frac{1}{4} \\ \frac{x}{2}&=\frac{1}{4}+\frac{2}{3} \\ \frac{x}{2}&=\frac{1\times 3}{4 \times 3}+\frac{2 \times 4}{3 \times 4} \\ \frac{x}{2}&=\frac{3}{12}+\frac{8}{12} \\ \frac{x}{2}&=\frac{11}{12} \end{align*} \] Finish the last step oin your own.

OpenStudy (anonymous):

yes but the answer is actually 1 5/6

OpenStudy (thomas5267):

Yes. It is \(1\frac{5}{6}\).